A plane can fly 1240 miles with a certain tail wind in 4 hours. When flying against a wind with the same velocity it can fly 1500 in 6 hours. What is the planes speed in still air?
Answer by askmemath(368)

Let the speed of the plane = S
Let speed of Wind = W
Now speed of plane per hour = 1240/4 = 310
Since the tail wind aids the plane S+W = 310
Similarly S-W = 1500/6 = 250
S+W = 310
S-W = 250
Adding both
2S = 560
S = 280
Therefore speed of plane = 280